The magnetic field in a region of space is specified by its x, y, and z components which are respectively Bx 0.138 T, By 0.162 T, and B 0.200 T. If a 30.0-cm wire carrying a current of 3.00 A is oriented along the z-axis with the current in the z direction, determine the magnitude and direction of the force acting on the wire. magnitude What are the components of the magnetic force? Can you determine the magnitude and direction of the magnetic force from the components of the force? N direction counterclockwise from the +X-axis
Answer #1

Given the magnetic field Bx = 0.138 T, By = 0.162 T, Bz = 0.200

length of the wire L = 0.3 m, curent in the -ve z direction is A

we know that the force F = I lB sin theta

           F = lI X B
cross product of I,B is

i j k

0 0 -3

0.138   0.162 0.200

i(0 + 3*0.162) -j (0+3*0.138)+k(0-0)

F = 0.3((0.486) i - 0.414 j + 0k)

F = 0.1458 i - 0.1242 j +0k


magnitude is F = sqrt((0.1458)^2+(-0.1242)^2) N = 0.1915287967905 N = 0.192 N


the direction is theta = arc tan (-0.1242/0.1458) = -40.426 degrees